# Resource Allocation via Backscatter-Aware Transmit Antenna Selection for Low-PAPR and Ultra-Reliable WSNs

### v1

@@ Abstract
This paper addresses the fundamental physical-layer conflict in hybrid Wireless Sensor Networks (WSNs) between high-throughput primary communication and the sensitive power-envelope requirements of passive sensors. We propose the Backscatter-Constrained Transmit Antenna Selection (BC-TAS) framework, a novel per-subcarrier selection scheme designed for multi-antenna illuminators operating within a Multi-Dimensional Orthogonal Frequency Division Multiplexing (MD-OFDM) architecture. Unlike conventional signal-to-noise ratio (SNR) centric selection, BC-TAS utilizes a multi-objective cost function to jointly optimize primary link reliability, stabilize the incident energy envelope at passive Surface Acoustic Wave (SAW) sensors, and provide surgical interference nulling toward coexisting victim nodes. To ensure robustness in low-power environments, we integrate a Kalman-based channel smoothing layer that maintains selection stability even under severe Channel State Information (CSI) estimation errors of up to 20%. Numerical evaluations using the non-linear Rapp power amplifier model and IEEE 802.11be dispersive channel models demonstrate that BC-TAS achieves a nearly three-order-of-magnitude reduction in outage probability and a $25\times$ improvement in network energy efficiency compared to standard MU-MIMO baselines. Furthermore, the proposed dual-envelope control mechanism achieves a $2.4$ dB reduction in the transmitter’s Peak-to-Average Power Ratio (PAPR), allowing for a $3$ dB reduction in required input back-off (IBO) while ensuring spectral mask compliance. These results confirm that BC-TAS effectively bridges the “Illuminator’s Gap,” providing active regulation of the RF envelope to ensure hardware linearity and robust sensing-communication coexistence in high-density, power-constrained environments.
@@ I Introduction
The rapid proliferation of Internet of Things (IoT) devices and their integration into large-scale Wireless Sensor Networks (WSNs) has intensified the demand for hybrid communication architectures. Modern industrial and medical environments now require a single active hub—the illuminator—to support high-throughput data transmission to legitimate users while simultaneously energizing and interrogating passive, backscatter-based sensors [Zhang2025_SR, Wu2024_WPCN]. However, this integration creates a fundamental physical-layer conflict: the wideband, high-power waveforms required for reliable primary communication often saturate the sensitive dynamic range of passive tags or cause prohibitive interference to coexisting victim receivers in the same spectral vicinity [Tan2025_Coexist].
@@
Conventional Multiple-Input Multiple-Output (MIMO) techniques, such as Minimum Mean Square Error (MMSE) precoding, are typically optimized for spectral efficiency. However, they are increasingly unsuitable for the next generation of energy-constrained Passive WSNs (PWSNs) due to three critical limitations. First, the requirement for multiple active RF chains imposes a hardware complexity and power consumption floor that is incompatible with green-WSN goals [Zhao2024_Green]. Second, the high Peak-to-Average Power Ratio (PAPR) inherent in multi-stream Orthogonal Frequency Division Multiplexing (OFDM) necessitates significant Power Amplifier (PA) back-off, leading to spectral regrowth and accelerated battery depletion [Park2024_PAPR]. Most importantly, current precoding strategies lack a mechanism to regulate the incident RF envelope perceived by remote passive devices [Wang2024_Sense]. For time-modulated sensors, such as Surface Acoustic Wave (SAW) tags, high-PAPR fluctuations act as a source of “envelope noise,” masking low-power acoustic reflections and degrading sensing fidelity [Liu2025_SAW].
@@
While diversity-oriented selection methods, such as Transmit Antenna Selection (TAS) and Multi-Dimensional OFDM (MD-OFDM), provide a low-complexity alternative with inherent PAPR benefits, they remain largely “sensing-blind” [Garcia2024_IM]. Existing selection literature focuses exclusively on maximizing desired-link Signal-to-Noise Ratio (SNR). Consequently, there is an urgent need for a framework that treats the spatial-frequency resource blocks of an OFDM symbol as a tool for multi-objective field shaping—simultaneously optimizing primary link reliability, stabilizing backscatter energy incidence, and surgically nulling interference toward victim nodes [Ibrahim2025_MultiObj].
@@
To address these gaps, we propose the Backscatter-Constrained Transmit Antenna Selection (BC-TAS) framework. This is the first per-subcarrier selection scheme designed to bridge the “Illuminator’s Gap” by actively managing the RF conditions perceived by passive devices. By shifting the selection logic from a purely SNR-centric approach to a multi-objective sensing-aware heuristic, BC-TAS transforms the illuminator to act as a network-wide controller of both signal quality and incident energy stability.
@@
The central innovation of this work is the development of a unified cost function that embeds three interacting channel components—the legitimate link, tag-side incidence, and victim-link interference—into a per-subcarrier decision rule. Furthermore, we leverage the sparsity of MD-OFDM to achieve a dual-regulation effect: minimizing transmitter-side PAPR to ensure PA linearity while simultaneously minimizing the Backscatter Crest Factor (BCF) to ensure tag-side envelope stability. This dual-regulation ensures that the system remains compliant with strict spectral masks, such as those defined in the maturing IEEE 802.11 Ambient Power Communication (AMP) frameworks, even when operating in highly dispersive environments [IEEE80211AMP_Update, IEEE80211bk_2025].
@@
  • BC-TAS Framework Design: We introduce a novel multi-objective per-subcarrier selection metric that explicitly balances primary link reliability against the stability of backscatter incidence and the suppression of victim-link interference.
  • Analytical Metric for Sensing Fidelity (BCF): We define and derive the BCF as a novel performance metric to quantify the temporal stability of the RF envelope at the tag. We provide a theoretical bound for BCF reduction under $N_{t}$-order selection diversity, establishing the first formal link between spatial resource allocation and sensing ”envelope noise.
  • Dual-Envelope Regulation: We establish a joint control mechanism that exploits sparse subcarrier activation to simultaneously reduce the transmitter’s PAPR and the tag’s incident BCF, providing the first quantitative link between illuminator antenna selection and backscatter sensing fidelity.
  • Low-Complexity Heuristic: We formulate the BC-TAS decision as a constrained optimization problem and develop a low-complexity greedy solution incorporating recursive Kalman-based channel smoothing to ensure selection stability under imperfect CSI. The algorithm scales linearly with the number of subcarriers, providing a 2.3× execution speedup over traditional MMSE detectors while maintaining robustness to rapid channel fluctuations.
  • Non-linear Hardware Validation: We perform a comprehensive evaluation using the Rapp PA model and IEEE 802.11n/be channel models, demonstrating over two orders of reduction in outage probability and a $25\times$ improvement in system energy efficiency compared to standard SISO and MIMO baselines.
The remainder of this paper is organized as follows. Section II reviews the state-of-the-art in TAS, backscatter sensing, and PAPR-aware transmission. Section III details the system model and the MOFS formulation, including the Kalman-based channel smoothing. Section IV presents the proposed greedy selection algorithm and derives its theoretical diversity and scaling properties. Section V provides extensive simulation results, including BER, BCF, and hardware efficiency metrics. Section VI synthesizes these findings with a theoretical validation of the observed gains and trade-offs. Finally, Section VII concludes the work, followed by the derivation of the selection penalty in Appendix A.
@@ II Related Work
II-A Transmit Antenna Selection (TAS)
TAS remains a cornerstone for low-complexity diversity in modern wireless architectures. Recent research by Kumar et al. (2025) demonstrates that TAS provides significant diversity gains in active RIS-assisted downlink NOMA systems, where the diversity order scales jointly with BS antennas and RIS elements [Kumar2025]. Similarly, Han et al. (2025) emphasize that TAS-based strategies, particularly under statistical Channel State Information (CSI), can effectively improve outage probability and sum throughput without the prohibitive RF-chain costs of full MIMO [Han2025]. However, these classical and 6G-era TAS frameworks primarily optimize internal communication metrics (SNR, Capacity, BER). There remains a critical research gap in applying TAS to manage external physical-layer constraints, such as the RF envelope of a backscatter tag or the interference profile of a sensitive victim node.
II-B OFDM PAPR Reduction and Hardware Linearity
@@
PAPR is the primary driver of PA non-linearity, leading to spectral regrowth and constellation distortion. Contemporary work in Integrated Sensing and Communication (ISAC) has highlighted that PAPR reduction is essential for improving transmitter efficiency and range under non-linear High Power Amplifier (HPA) constraints [Huang2022]. While probabilistic methods like Partial Transmit Sequences (PTS) have been refined to embed side information implicitly to reduce overhead [Goel2022], they still introduce signal processing latency. From a system-level perspective, Komala (2024) demonstrates that inefficient transmission directly translates to reduced network longevity in WSNs [Komala2024]. By mapping these findings to the Rapp model, it becomes evident that an inherently low-PAPR waveform, rather than an algorithmic mitigation layer, is vital for long-lived, energy-efficient Wireless Power and Data Transfer (WPDT).
II-C Index Modulation and MD-OFDM
Index Modulation (IM) and its variants, such as MD-OFDM, have emerged as core 6G enablers for massive connectivity [Li2023]. Unlike classical OFDM, IM exploits the indices of subcarriers or antennas to embed information, providing a sparse activation pattern that can lead to lower PAPR and reduced inter-channel interference [Mgobhozi2023]. Recent advances in permutational IM have even enabled low-complexity, non-coherent detection that is robust under antenna correlation [VilaInsa2025]. However, most IM literature focuses on maximizing spectral efficiency (SE) or reliability (BER). Even when PAPR is addressed, as in multilevel dither signal designs for OFDM-IM [Kim2018], the focus is purely on internal signal integrity. Our work bridges this gap by treating the MD-OFDM index selection as a first-class design variable for shaping the incident energy at a backscatter sensor.
II-D Backscatter Communications and Symbiotic Radio
@@
The evolution of backscatter has moved toward multi-antenna and broadband configurations to improve link reliability in dispersive channels [Zhang2017, Chen2024]. Optimization efforts have traditionally been split: either optimizing tag reflection coefficients for energy-rate trade-offs [Goay2024] or utilizing deep reinforcement learning for beamforming in bistatic setups [Zargari2024]. Recent symbiotic radio research has proposed strategic subcarrier nulling to eliminate direct-link interference in OFDM backscatter [Janjua2025]. Despite these advances, the ”Illuminator’s Gap” persists; few studies investigate using the transmitter’s resource allocation (like BC-TAS) to actively shape the temporal stability of the RF envelope, which is critical for time-modulated sensors like SAW tags.
II-E Coexistence in Passive SAW-based Sensor Networks
As standardization bodies like IEEE 802.11 extend positioning to 320 MHz channels [IEEE80211bk], the need for interference-resilient signaling becomes paramount. Passive SAW sensors represent a unique challenge in this landscape, as they operate without batteries by converting incident RF energy into time-delayed acoustic reflections [Chen2024, Zhang2017]. Unlike active IoT devices, PWSNs are highly susceptible to:
  • RF Envelope Instability: High PAPR signals create power fluctuations that mask the low-power acoustic reflections [Huang2022, Kim2018].
  • Spectral Competition: While frequency-domain nulling has been proposed [Liu2025, Janjua2025], these methods often ignore the hardware non-linearities of the illuminator.
Industry consensus in the IEEE 802.11 AMP group highlights a critical gap: the absence of selection-based mechanisms that utilize the transmitter’s spatial degrees of freedom to optimize the energy envelope at the tag while protecting primary link reliability. The performance of selection-based architectures in industrial settings is inextricably linked to the underlying radio propagation environment. Prior work has introduced Variational Autoencoder (VAE) based models to predict signal-to-interference-plus-noise ratio (SINR) heatmaps within automated Industry 4.0 warehouses, demonstrating how interacting objects within smart infrastructure significantly impact 5G propagation [GuliaVAE2025]. Furthermore, detailed evaluations of 60 GHz millimeter-wave connectivity in these environments reveal that non-line-of-sight (nLOS) paths created by metallic shelving and autonomous handling agents introduce complex multipath fading effects [Gulia60GHz2023]. To manage these dynamics without high computational costs, stabilized symbolic regression has been utilized to derive compact, interpretable expressions for block error rate (BLER) prediction, offering neural-network-level accuracy suitable for real-time embedded deployment [GuliaSymbolic2024]. These studies underscore the necessity for low-complexity, environment-aware selection mechanisms like BC-TAS that can adapt to the complex spatial profiles of automated warehouses.
Recognizing these coexistence challenges, the industrial shift toward sustainable IoT is evidenced by the formation of the IEEE 802.11 AMP Task Group (TGbp). According to the project authorization scope, the group aims to enable battery-free station (STA) operation through energy harvesting and ultra-low-complexity signaling [IEEE80211AMP_PAR]. A primary challenge identified in the group’s functional requirements is the coexistence of these passive-like devices with legacy high-throughput STAs [IEEE80211AMP_Update]. While the task group explores physical-layer modifications for 2.4 GHz and sub-1 GHz bands, current proposals primarily focus on frequency-domain resource allocation. Our work, BC-TAS, complements these efforts by introducing a spatial-domain selection mechanism that addresses the hardware non-linearities and peak-envelope requirements that are not yet fully specified in the maturing AMP framework.
@@ IV-C Optimality Gap and Heuristic Justification
V Simulation and Results
This section details the simulation environment, defines the key parameters used, and presents the performance metrics of the proposed Backscattered Constrained Transmit Antenna Selection Multi-Dimensional Orthogonal Frequency Division Multiplexing (BC-TAS MD-OFDM) model compared to baseline schemes.
V-A Simulation Setup
The system performance was evaluated using Monte Carlo simulations. The core parameters are summarized in Table I. To ensure relevance to emerging standards, the channel environment is modeled based on the IEEE 802.11be (Wi-Fi 7) TGbe indoor specifications, utilizing Multipath Fading Models (TGn A–F) to represent varying degrees of frequency selectivity in warehouse environments.
To characterize hardware-induced distortions, we incorporate a non-linear Rapp PA model at the transmitter. The AM/AM conversion follows the smoothness factor $p=2$, which represents a typical Solid State Power Amplifier (SSPA). This model is critical for evaluating the impact of high-PAPR and high-BCF waveforms on spectral regrowth and Legitimate-link Bit Error Rate (BER).
Simulations were performed on a high-performance computing node equipped with an NVIDIA Tesla V100-PCIE-16GB GPU and Intel Xeon processors, utilizing the NVIDIA CUDA 12.3 compiler driver. Execution times reported in Table II represent the total duration required to process $10^{7}$ bits across an SNR range of 0 to 28 dB.

### v2

@@ Abstract
This paper addresses the fundamental physical-layer conflict in hybrid Wireless Sensor Networks (WSNs) between high-throughput primary communication and the sensitive power-envelope requirements of passive sensors. We propose the Backscatter-Constrained Transmit Antenna Selection (BC-TAS) framework, a novel per-subcarrier selection scheme designed for multi-antenna illuminators operating within a Multi-Dimensional Orthogonal Frequency Division Multiplexing (MD-OFDM) architecture. Unlike conventional signal-to-noise ratio (SNR) centric selection, BC-TAS utilizes a multi-objective cost function to jointly optimize primary link reliability, stabilize the incident energy envelope at passive Surface Acoustic Wave (SAW) sensors, and provide surgical interference nulling toward coexisting victim nodes. To ensure robustness in low-power environments, we integrate a Kalman-based channel smoothing layer that maintains selection stability even under severe Channel State Information (CSI) estimation errors of up to 20%. Numerical evaluations using the non-linear Rapp power amplifier model and IEEE 802.11be dispersive channel models demonstrate that BC-TAS achieves a nearly three-order-of-magnitude reduction in outage probability and a $25\times$ improvement in network energy efficiency compared to standard Multi-User Multiple Input, Multiple Output (MU-MIMO) baselines. Furthermore, the proposed dual-envelope control mechanism achieves a $2.4$ dB reduction in the transmitter’s Peak-to-Average Power Ratio (PAPR), allowing for a $3$ dB reduction in required input back-off (IBO) while ensuring spectral mask compliance. These results confirm that BC-TAS effectively bridges the “Illuminator’s Gap,” providing active regulation of the RF envelope to ensure hardware linearity and robust sensing-communication coexistence in high-density, power-constrained environments.
@@ I Introduction
The proliferation of Internet of Things (IoT) devices in large-scale Wireless Sensor Networks (WSNs) has intensified the demand for hybrid communication architectures. Modern industrial and medical environments require a single active hub—the illuminator—to support high-throughput data transmission while simultaneously energizing and interrogating passive, backscatter-based sensors [Zhang2025_SR, Wu2024_WPCN]. However, this integration creates a fundamental physical-layer conflict: the wideband, high-power waveforms required for primary communication often saturate the sensitive dynamic range of passive tags or cause prohibitive interference to coexisting victim receivers [Tan2025_Coexist].
@@
Conventional Multiple-Input Multiple-Output (MIMO) techniques, such as Minimum Mean Square Error (MMSE) precoding, are increasingly unsuitable for energy-constrained Passive WSNs (PWSNs) due to three critical limitations. First, multiple active RF chains impose a hardware complexity and power floor incompatible with green-WSN goals [Zhao2024_Green]. Second, the high Peak-to-Average Power Ratio (PAPR) of multi-stream OFDM necessitates significant Power Amplifier (PA) back-off, leading to spectral regrowth and accelerated battery depletion [Park2024_PAPR]. Most importantly, current strategies lack a mechanism to regulate the incident RF envelope perceived by remote passive devices [Wang2024_Sense]. For time-modulated sensors, such as Surface Acoustic Wave (SAW) tags, high-PAPR fluctuations act as “envelope noise,” masking low-power acoustic reflections and degrading sensing fidelity [Liu2025_SAW].
@@
While selection methods like Transmit Antenna Selection (TAS) and Multi-Dimensional OFDM (MD-OFDM) provide low-complexity alternatives with inherent PAPR benefits, they remain largely “sensing-blind” [Garcia2024_IM]. Existing literature focuses almost exclusively on maximizing desired-link Signal-to-Noise Ratio (SNR). Consequently, there is an urgent need for a framework that treats spatial-frequency resource blocks as tools for multi-objective field shaping—simultaneously optimizing primary link reliability, stabilizing backscatter energy incidence, and surgically nulling interference toward victim nodes [Ibrahim2025_MultiObj].
@@
To address these gaps, we propose the Backscatter-Constrained Transmit Antenna Selection (BC-TAS) framework. To the best of our knowledge, this is the first per-subcarrier selection scheme designed to bridge the “Illuminator’s Gap” by actively managing the RF conditions perceived by passive devices. By shifting selection logic from an SNR-centric approach to a multi-objective sensing-aware heuristic, BC-TAS transforms the illuminator into a network-wide controller of both signal quality and incident energy stability.
@@
The central innovation is a unified cost function embedding three interacting channel components—the legitimate link, tag-side incidence, and victim interference—into a per-subcarrier decision rule. We leverage the sparsity of MD-OFDM to achieve a dual-regulation effect: minimizing transmitter-side PAPR for PA linearity while simultaneously minimizing the Backscatter Crest Factor (BCF) for tag-side envelope stability. This ensures compliance with strict spectral masks, such as the IEEE 802.11bp Ambient Power Communication (AMP) frameworks, even in highly dispersive environments [IEEE80211AMP_Update].
@@
  • BC-TAS Framework Design: A novel multi-objective per-subcarrier selection metric that balances primary link reliability against backscatter incidence stability and victim interference suppression.
  • Analytical Metric for Sensing Fidelity BCF: We define the BCF to quantify the temporal stability of the RF envelope at the tag and provide a theoretical bound for BCF reduction under $N_{t}$-order selection diversity.
  • Dual-Envelope Regulation: A joint control mechanism exploiting sparse subcarrier activation to simultaneously reduce transmitter PAPR and tag incident BCF, linking antenna selection directly to sensing fidelity.
[…]
  • Hardware Aware Simulation: Comprehensive evaluation using Rapp PA and IEEE 802.11 TGn models, demonstrating two orders of magnitude reduction in outage probability and a $25\times$ improvement in energy efficiency.
The remainder of this paper is organized as follows. Section II reviews the state-of-the-art. Section III details the system model and Multi-Objective Field Shaping (MOFS) formulation. Section IV presents the greedy selection algorithm and its theoretical properties. Section V provides extensive simulation results and synthesizes these with a theoretical validation. Section VI concludes the work.
@@ II Related Work
II-A Transmit Antenna Selection (TAS) and Index Modulation
TAS remains a cornerstone for low-complexity diversity. Recent research demonstrates significant gains in RIS-assisted NOMA systems and 6G-era architectures where diversity scales with antenna count without prohibitive RF-chain costs [Kumar2025, Han2025]. Parallelly, Index Modulation (IM) and MD-OFDM have emerged as massive connectivity enablers by exploiting subcarrier indices to embed information [Li2023]. While IM provides sparse activation and reduced interference [Mgobhozi2023, VilaInsa2025], existing TAS and IM literature primarily optimizes internal metrics like SNR and Bit Error Rate (BER). A critical gap remains in utilizing these selection variables to manage external constraints, such as the RF envelope of a backscatter tag.
II-B PAPR, Hardware Linearity, and Energy Efficiency
@@
PAPR is the primary driver of PA non-linearity, leading to spectral regrowth and constellation distortion. In Integrated Sensing and Communication (ISAC), PAPR reduction is vital for efficiency under non-linear High Power Amplifier (HPA) constraints [Huang2022]. While probabilistic methods like Partial Transmit Sequences (PTS) reduce PAPR, they introduce processing latency that impacts WSN longevity [Goel2022, Komala2024]. Mapping these constraints to the Rapp model suggests that inherently low-PAPR waveforms are superior to algorithmic mitigation for energy-efficient Wireless Power and Data Transfer (WPDT).
II-C Backscatter Communications and Symbiotic Radio
@@
Backscatter evolution has moved toward multi-antenna and broadband configurations to enhance link reliability [Chen2024]. Optimization efforts often split between tag-side reflection coefficients and bistatic beamforming [Goay2024, Zargari2024]. While symbiotic radio research proposes subcarrier nulling to eliminate direct-link interference [Janjua2025], few studies investigate using transmitter resource allocation to shape the temporal stability of the RF envelope—a factor critical for time-modulated SAW tags.
II-D Coexistence in Passive SAW-based Sensor Networks
As IEEE 802.11 positioning extends to 320 MHz channels [IEEE80211bk], interference-resilient signaling is paramount. Passive SAW sensors are particularly susceptible to RF envelope instability, where high PAPR masks low-power acoustic reflections [Huang2022, Kim2018]. While frequency-domain nulling has been proposed [Liu2025, Janjua2025], these methods often ignore hardware non-linearities. Industry consensus within the IEEE 802.11 AMP group (TGbp) highlights the lack of selection-based mechanisms to optimize tag energy envelopes while protecting primary link reliability [IEEE80211AMP_Update].
Selection performance in industrial settings is inextricably linked to the propagation environment. Prior work utilizing VAE-based SINR heatmaps and 60 GHz connectivity analysis in automated warehouses reveals complex multipath fading caused by metallic infrastructure [GuliaVAE2025, Gulia60GHz2023]. To manage these dynamics, interpretable symbolic regression has been used for real-time BLER prediction [GuliaSymbolic2024]. These studies underscore the necessity for environment-aware mechanisms like BC-TAS. Our framework complements current TGbp efforts by introducing a spatial-domain selection mechanism that addresses hardware non-linearities and peak-envelope requirements not yet fully specified in the maturing AMP framework [IEEE80211AMP_PAR].
@@ IV-C Optimality Gap and Heuristic Justification
V Computational Complexity Analysis
The feasibility of BC-TAS for real-time industrial deployment is rooted in its linear scaling properties. Unlike conventional MMSE precoding, which requires matrix inversions of the form $\mathcal{O}(K^{3})$ (where $K$ is the number of users/streams), the proposed greedy selection algorithm operates on a per-subcarrier basis with a complexity of $\mathcal{O}(N_{SC}\cdot N_{t})$.
Specifically, for each subcarrier $n\in\{1,\dots,N_{SC}\}$, the metric $M(n,i)$ is evaluated $N_{t}$ times. The inclusion of the scalar Kalman Filter adds a constant overhead of $\mathcal{O}(1)$ per subcarrier, as it avoids high-dimensional state-space computations. As a result, for a standard 1024-subcarrier system with 4 transmit antennas, BC-TAS reduces the total floating-point operations (FLOPs) by approximately $64\%$ compared to full-complexity MIMO baselines. This reduction directly translates to the $2.3\times$ execution speedup reported in our results, enabling the illuminator to adapt to the fast-fading environments of automated warehouses.
VI Simulation and Results
This section evaluates the proposed BC-TAS MD-OFDM framework against state-of-the-art (SOTA) baselines across indoor dispersive environments.
VI-A Simulation Setup
System performance is evaluated via Monte Carlo simulations using the parameters in Table I. To ensure standard relevance, the channel is modeled on IEEE 802.11 TGn specifications, utilizing Multipath Fading Models (TGn A–F) to represent varying frequency selectivity in warehouse environments. Hardware distortions are characterized by a non-linear Rapp PA model ($p=2$) at the transmitter to evaluate the impact of high-PAPR/BCF waveforms on spectral regrowth and Legitimate-link Bit Error Rate (BER).
Simulations were conducted on a high-performance node (NVIDIA Tesla V100 GPU, Intel Xeon processors) with the CUDA 12.3 driver. Execution times in Table II reflect the processing of $10^{7}$ bits across an SNR range of 0–28 dB.